# A notion of fractality for a class of states and noncommutative relative distance zeta functional

### v3

@@ 4 Examples
(A commutative example.) Let $d=1$, $n=1$, and write $H=\mathbb{R}$, and hence $\mathfrak{sa}\,\mathcal{L}(H)=\mathbb{R}$. Let $\mu=\mu_{\mathcal{C}}(dy)$, $\nu=\eta(x)dx$ for a compactly supported $\eta$ that covers [1,2][-1,2]. We may compare this with the results of [41] about the Cantor set and Centor graph, and their associate relative fractal drums.
@@
We note that in fact, with the above procedure and a similar argument as in [11], it is expected that we shall obtain a general functional equation using Lemmas 3.2 and 3.5 whenever a general self-similar structure arises. Similar functional equations satisfied by the geometric, fractal and distance zeta function with rescaling were previously also introduced and used earlier in different context, including [12, 13, 14, 15, 16, 21, 25, 26, 27, 34, 35, 36, 37, 42, 43, 45]. This is a direction worth pursuing in our next work, especially rigorously establishing concrete criteria so as to guarantee an associated scaling zeta function similar to Definition 4.2 in [11] alone may recover all the poles of our relative distance zeta functional.

### v4

@@ 4 Examples
(A commutative example.) Let $d=1$, $n=1$, and write $H=\mathbb{R}$, and hence $\mathfrak{sa}\,\mathcal{L}(H)=\mathbb{R}$. Let $\mu=\mu_{\mathcal{C}}(dy)$, $\nu=\eta(x)dx$ for a compactly supported $\eta$ that covers [1,2][-1,2]. We may compare this with the results of [42] about the Cantor set ([42], Section 1.1, Example 1.1.2) and the graph of the Cantor function ([42], Section 5.5.4, Example 5.5.14), and their associate relative fractal drums.
@@
We note that in fact, with the above procedure and a similar argument as in [11], it is expected that we shall obtain a general functional equation using Lemmas 3.2 and 3.5 whenever a general self-similar structure arises. Similar functional equations satisfied by the geometric, fractal and distance zeta function with rescaling were previously also introduced and used earlier in different context, including [12, 13, 14, 15, 16, 21, 25, 26, 27, 34, 35, 36, 37, 42, 43, 45]. This is a direction worth pursuing in our next work, especially rigorously establishing concrete criteria so as to guarantee an associated scaling zeta function similar to Definition 4.2 in [11] and in [35, 36, 42] alone may recover all the poles of our relative distance zeta functional.